OpenStudy (anonymous):

.Mary is 14 years younger than her sister Jane. In four years, Jane will be 3 times older than Mary. How old is Jane now?

OpenStudy (anonymous):

@Calcmathlete please could you help

OpenStudy (anonymous):

@twiggyc That is incorrect.

OpenStudy (anonymous):

Here's a system of equations. \[M = J - 14\]\[J + 4 = 3(M + 4)\]Can you trying figuring it out?

OpenStudy (anonymous):

You can easily substitute in what M is.

OpenStudy (anonymous):

The reasonb yours didn't work is because you forgot to account for Mary's age when 4 years has passed.